# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v1

@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
Figure 2: The experimentally measured and theoretically predicted values for $R_{D_{0}^{*}}$. VMD__\__ 2021 is the result obtained by VMD approach where $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff ; Fornal:2017msy . The result of VMD__\__ 2023 denoted in blue is calculated by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and that of VMD__\__ 2023 denoted in black is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. The SM prediction is denoted as SM__\__ Photon.
@@
We observe that the above adjustments of $\varepsilon_{c}$ and $\varepsilon_{s}$ naturally affect the decays of Charmonium ($c\bar{c}$) and $\phi$ ($s\bar{s}$) mesons. Hence the determination of these coupling parameters should involve the data of such decay modes.
@@ III X17 couplings to charm and strange quarks
The VMD assumption has been phenomenologically successful in explaining TFF behaviors in similar decays. The effective pole mass refers to the mass of the vector meson resonance that is close to the energy scale of the decaying particle. In the analysis of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\Lambda_{\psi(2S)\eta_{c}}$ corresponds to the mass of a higher excited state, $m_{\psi(3770)}=(3773.7\pm 0.4)\,\mathrm{MeV/c^{2}}$ Gu:2019qwo . This framework has provided a coherent explanation for the experimental observations in the Charmonium decays. From Eq. (7), the ratio $R_{\psi(2S)}^{X}$ between the decay rate of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$ decay rate and that of $\psi(2S)\rightarrow\eta_{c}\gamma$ is given by
@@
This is an approximated result in the limit $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. The measurement by BESIII gives ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}\right)=$ $\left(3.77\pm 0.40_{\text{stat }}\pm 0.18_{\text{syst }}\right)\times 10^{-5}$ BESIII:2022jde . Utilizing the branching ratio ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}\gamma\right)=(3.6\pm 0.5)\times 10^{-3}$ ParticleDataGroup:2024cfk , one obtains the ratio $R^{\rm data}_{\psi(2S)}=(1.1\pm 0.2)\times 10^{-2}$ as reported in BESIII:2022jde as well. The theoretical prediction for the photon-mediated decay rate given by the VMD model is $R_{\psi(2S)}^{\gamma}=8.9\times 10^{-3}$. The tiny difference between $R_{\psi(2S)}^{\gamma}$ and $R^{\rm data}_{\psi(2S)}$ gives a strong constraint to $R_{\psi(2S)}^{X}$, i.e., the coupling parameter $\varepsilon_{c}$.
@@
Similar to the case of $\psi(2S)$ decay, we obtain this result with the approximations $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\phi\eta\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. Given the measurements on $R_{\phi}$ from three separate experiments and the theoretical prediction, $R_{\phi}^{\gamma}=8.30\times 10^{-3}$, we are able to determine the allowed range for $R_{\phi}^{X}$.
@@
Figure 3: The allowed ranges for (a) $\left(\varepsilon_{c},\varepsilon_{s}\right)$ and (b) $\left(\varepsilon_{c},\varepsilon_{u}\right)$ extracted by the data of $D^{*+}_{s}$, $D^{*0}$, $\psi(2S)$ and $\phi$ decays.
The $1\sigma$ parameter range allowed by the ratio between $\Gamma(\psi(2S)\to\eta_{c}e^{+}e^{-})$ BESIII:2022jde and $\Gamma(\psi(2S)\to\eta_{c}\gamma)$ is denoted by the pink area in Fig. 3. The left panel depicts the allowed range for the $\left(\varepsilon_{c},\varepsilon_{s}\right)$ pair while the right panel depicts that for $\left(\varepsilon_{c},\varepsilon_{u}\right)$. Furthermore, the left panel of Fig. 3 also shows $1\sigma$ ranges for $\varepsilon_{s}$ from the data of SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk on $\phi\to\eta e^{+}e^{-}$, denoted by yellow, gray, and purple areas, respectively. The blue band is the allowed range for $\left(\varepsilon_{c},\varepsilon_{s}\right)$ by the data of $D^{*+}_{s}\to D^{+}_{s}e^{+}e^{-}$ CLEO:2011mla . On the right panel of Fig. 3, the $1\sigma$ allowed range for $\left(\varepsilon_{c},\varepsilon_{u}\right)$ by the data of $D^{*0}\to D^{0}e^{+}e^{-}$ is denoted by the green areas. Finally the extracted range for $|\varepsilon_{u}|$ from Refs. Feng:2016jff and  Denton:2023gat are denoted by black dashed curves and red band, respectively.
@@
$\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ $(11\pm 1)\times 10^{-3}$ $6.7\times 10^{-3}$
@@
$\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ $(11\pm 2)\times 10^{-3}$ $8.9\times 10^{-3}$
@@
$\chi^{2}=\sum_{i=1}^{4}\frac{\left(R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u}\right)-R_{i}^{\rm ob}\right)^{2}}{\sigma_{i}^{2}}+\frac{\left(\lambda-\lambda_{\rm best\,fit}\right)^{2}}{\sigma_{\lambda}^{2}},$ (18)
where $R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u}\right)$ denotes the theoretically predicted value for the $i$-th decay ratio, which depends on the parameters $\varepsilon_{c}$, $\varepsilon_{s}$, and $\varepsilon_{u}$. $R_{i}^{\rm ob}$ refers to the observed value for the $i$-th decay ratio with $\sigma_{i}$ the corresponding uncertainty for the measurement. The parameter $\lambda$ determines the coupling strength of the heavy meson to light vector mesons, i.e., it appears in the coupling vertex $\lambda H^{*}HV$. From Eq. (9), we have $\lambda_{\rm best\,fit}=-0.289$ GeV-1 and $\sigma_{\lambda}=0.016$ GeV-1. Our fitting uses the measured decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons as outlined in Table 2.
@@
The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=3\times 10^{-4}$ and $\left|\varepsilon_{s}\right|=4\times 10^{-5}$ while there are two best-fit solutions for $\left|\varepsilon_{u}\right|$, which are $4.7\times 10^{-2}$ and $5.0\times 10^{-2}$, respectively. The fitting results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. Up to the $3\sigma$ range, the value of $\left|\varepsilon_{u}\right|$ is of the order $10^{-2}$, which is notably larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ Denton:2023gat . In other words, the value of $\left|\varepsilon_{u}\right|$ extracted from the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons is in a serious tension with that extracted by ATOMKI measurements. The significant increase of $\left|\varepsilon_{u}\right|$ in this case can be attributed to the anomalously large branching ratio of $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ measured by by BESIII BESIII:2021vyq as illustrated by Fig. 2 and the right panel of Fig. 3. We remark that only the decay of $D^{*0}$ among those of $D_{s}^{*+}$, $D^{*0}$, $\psi(2S)$, and $\phi$ mesons involve X17 couplings to the first generation of quarks, i.e., $\varepsilon_{u}$ specifically.
@@ IV Conclusions
The confirmation of couplings between X17 boson and quarks remains a topic of interest to date. The hypothesis of X17 boson has continuously received attentions by the data of anomalous 8Be, 4He, and 12C decays over time. The accumulation of data has shifted the most favored coupling strengths from $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ Feng:2016jff to $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ without isospin effects, and $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ with both isospin mixing and breaking effects considered Denton:2023gat . The persistent decay anomalies outlined above have motivated the testing of X17 hypothesis in the decays of heavy-flavor mesons such as $D^{*}$ and $B^{*}$ mesons with the assumption $\varepsilon_{b}=\varepsilon_{s}=\varepsilon_{d}$ and $\varepsilon_{c}=\varepsilon_{u}$ Castro:2021gdf . We have revisited this study by focusing on the comparisons of theoretical predictions (with X17 boson mediated effects included) to decay ratios $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ with experimental data CLEO:2011mla ; BESIII:2021vyq . For the case of $R_{D_{s}^{*+}}$, we have found that only the updated values of $\varepsilon_{u,d}$ are consistent with the measurement while the earlier extracted parameter values $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ predict a $R_{D_{s}^{*+}}$ much greater than the measurement (see Fig. 1). Our finding corrects an error by Ref. Castro:2021gdf which concluded that the prediction by $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ is consistent with the measurement.
@@
To accommodate the measured $R_{D^{*0}}$, one requires a much larger $\left|\varepsilon_{u}\right|$ under the assumption $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, raising the magnitude of $\varepsilon_{c}$ would affect $R_{D_{s}^{*+}}$ unless the magnitude of $\varepsilon_{s}$ is increased as well since $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ as indicated by Fig. 3 (left panel). Since $\varepsilon_{s}$ already departs from $\varepsilon_{d}$ in this case, we then also treat $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent parameters in our analysis. To determine the favored ranges for $\varepsilon_{u}$, $\varepsilon_{c}$ and $\varepsilon_{s}$, we have performed a fitting to the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}$, and $R_{\phi}$. The results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=3\times 10^{-4}$, $\left|\varepsilon_{s}\right|=4\times 10^{-5}$, and two solutions for $\left|\varepsilon_{u}\right|$, which are $4.7\times 10^{-2}$ and $5.0\times 10^{-2}$, respectively.
@@
It is noteworthy that the value of $\left|\varepsilon_{u}\right|$ remains of the order $10^{-2}$ up to the $3\sigma$ range. Remarkably, this value is much larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|=(0.5-0.9)\times 10^{-3}$ Denton:2023gat . Such a tension is caused by the large deviation of the measured $R_{D^{*0}}$ from the Standard Model prediction as shown in Table 2. This is also reflected by the right panel of Fig. 3, where both green bands significantly deviate from the line $\varepsilon_{u}=0$. In conclusion, we have studied effects of the proposed X17 boson to the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons, which involve X17 boson couplings to both the first and second generations of quarks. The above-mentioned tension on the value of $\varepsilon_{u}$ deserves further scrutiny. We note that the decay $D^{*\pm}\to D^{\pm}e^{+}e^{-}$ remains to be measured. Furthermore the huge deviation of the measured $D^{*0}\to D^{0}e^{+}e^{-}$ branching ratio to its Standard Model predicted value awaits confirmations from updated measurements.

### v2

@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
Figure 2: The experimentally measured and theoretically predicted values for $R_{D_{0}^{*}}$. VMDa is the result obtained by VMD approach where $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff ; Fornal:2017msy . The result of VMDb denoted in blue is calculated by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and that of VMDb denoted in black is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. The SM prediction is denoted as SM__\__ Photon.
@@
We observe that the above adjustments of $\varepsilon_{c}$ and $\varepsilon_{s}$ naturally affect the decays of Charmonium ($c\bar{c}$) and $\phi$ ($s\bar{s}$) mesons. Hence the determination of these coupling parameters should involve the data of these decay channels.
@@ III X17 couplings to charm and strange quarks
The VMD assumption has been phenomenologically successful in explaining TFF behaviors in similar decays. The effective pole mass refers to the mass of the vector meson resonance that is close to the energy scale of the decaying particle. In the analysis of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\Lambda_{\psi(2S)\eta_{c}}$ corresponds to the mass of a higher excited state, $m_{\psi(3770)}=(3773.7\pm 0.4)\,\mathrm{MeV/c^{2}}$ Gu:2019qwo . This framework has provided a coherent explanation for the experimental observations in the Charmonium decays. From Eq. (7), the ratio $R_{\psi(2S)}^{X}$ between the decay rate of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$ and that of $\psi(2S)\rightarrow\eta_{c}\gamma$ is given by
@@
This is an approximated result in the limit $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. The measurement by BESIII gives ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}\right)=$ $\left(3.77\pm 0.40_{\text{stat }}\pm 0.18_{\text{syst }}\right)\times 10^{-5}$ BESIII:2022jde . Utilizing the branching ratio ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}\gamma\right)=(3.6\pm 0.5)\times 10^{-3}$ ParticleDataGroup:2024cfk , one obtains the ratio $R^{\rm exp}_{\psi(2S)}=(1.1\pm 0.3)\times 10^{-2}$ as reported in BESIII:2022jde . The theoretical prediction for the photon-mediated decay rate given by the VMD model is $R_{\psi(2S)}^{\gamma}=8.9\times 10^{-3}$. The tiny difference between $R_{\psi(2S)}^{\gamma}$ and $R^{\rm exp}_{\psi(2S)}$ strongly constrain $R_{\psi(2S)}^{X}$, i.e., the coupling parameter $\varepsilon_{c}$.
@@
Similar to the case of $\psi(2S)$ decay, we obtain the above relation with the approximations $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\phi\eta\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. Given the measurements on $R_{\phi}$ from three separate experiments and the theoretical prediction, $R_{\phi}^{\gamma}=8.30\times 10^{-3}$, we are able to determine the allowed range for $R_{\phi}^{X}$.
@@
Figure 3: The allowed ranges for (a) $\left(\varepsilon_{c},\varepsilon_{s}\right)$ and (b) $\left(\varepsilon_{c},\varepsilon_{u}\right)$ extracted by the data of $D^{*+}_{s}$, $D^{*0}$, $\psi(2S)$ and $\phi$ decays. The values of $\varepsilon_{u}$ extracted by Refs. Feng:2016jff ; Fornal:2017msy and Ref. Denton:2023gat are also shown.
The $1\sigma$ parameter range allowed by the ratio between $\Gamma(\psi(2S)\to\eta_{c}e^{+}e^{-})$ BESIII:2022jde and $\Gamma(\psi(2S)\to\eta_{c}\gamma)$ is denoted by the pink area in Fig. 3. The left panel depicts the allowed range for the $\left(\varepsilon_{c},\varepsilon_{s}\right)$ pair while the right panel depicts that for $\left(\varepsilon_{c},\varepsilon_{u}\right)$. Furthermore, the left panel of Fig. 3 also shows $1\sigma$ ranges for $\varepsilon_{s}$ from the data of SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk on $\phi\to\eta e^{+}e^{-}$, denoted by yellow, gray, and purple areas, respectively. The blue band is the allowed range for $\left(\varepsilon_{c},\varepsilon_{s}\right)$ by the data of $D^{*+}_{s}\to D^{+}_{s}e^{+}e^{-}$ CLEO:2011mla . On the right panel of Fig. 3, the $1\sigma$ allowed range for $\left(\varepsilon_{c},\varepsilon_{u}\right)$ by the data of $D^{*0}\to D^{0}e^{+}e^{-}$ is denoted by the green areas. The parameter ranges for various $D^{*}$ meson decays are determined with $\lambda=-0.289\ {\rm GeV}^{-1}$. Finally the extracted values for $\varepsilon_{u}$ from Refs. Feng:2016jff ; Fornal:2017msy and Ref. Denton:2023gat are denoted by black dashed curves and red band, respectively.
@@
$\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ $(11.08\pm 0.9)\times 10^{-3}$ $6.7\times 10^{-3}$
@@
$\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ $(11\pm 3)\times 10^{-3}$ $8.9\times 10^{-3}$
@@
$\chi^{2}=\sum_{i=1}^{4}\frac{\left(R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u},\lambda\right)-R_{i}^{\rm ob}\right)^{2}}{\sigma_{i}^{2}}+\frac{\left(\lambda-\lambda_{\rm best\,fit}\right)^{2}}{\sigma_{\lambda}^{2}},$ (22)
where $R_{i}^{\rm th}$ denotes the theoretically predicted value for the $i$-th decay ratio, which depends on the parameters $\varepsilon_{c}$, $\varepsilon_{s}$, $\varepsilon_{u}$, and $\lambda$ (for $i=D^{*0},\ D^{*+}_{s}$). $R_{i}^{\rm ob}$ refers to the observed value for the $i$-th decay ratio with $\sigma_{i}$ the corresponding uncertainty for the measurement. The parameter $\lambda$ is defined in Eq. (11) with its range given by Eq. (13). We hence take $\lambda_{\rm best\,fit}=-0.289$ GeV-1 and $\sigma_{\lambda}=0.016$ GeV-1. Our fitting uses the measured decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons as outlined in Table 2.
@@
The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=6.4\times 10^{-3}$ and $\left|\varepsilon_{s}\right|=2.0\times 10^{-3}$ while there are two best-fit solutions for $\left|\varepsilon_{u}\right|$, which are $6.0\times 10^{-2}$ and $6.5\times 10^{-2}$, respectively. We have $\chi^{2}_{\rm min}=0.46$ for the fitting. The fitting results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. Up to the $3\sigma$ range, the value of $\left|\varepsilon_{u}\right|$ is of the order $10^{-2}$, which is notably larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|\simeq(0.5-0.9)\times 10^{-3}$ Denton:2023gat . In other words, the value of $\left|\varepsilon_{u}\right|$ extracted from the decays of $D^{*}$, $\psi(2S)$, and $\phi$ mesons is in a serious tension with that extracted by ATOMKI measurements. The significant increase of $\left|\varepsilon_{u}\right|$ in this case can be attributed to the anomalously large branching ratio of $D^{*0}\rightarrow D^{0}e^{+}e^{-}$ measured by by BESIII BESIII:2021vyq as illustrated by Fig. 2 and the right panel of Fig. 3. We remark that only the decay of $D^{*0}$ among our considered decay modes involve X17 couplings to the first generation of quarks.
@@ IV Conclusions
The confirmation of couplings between X17 boson and quarks remains a topic of interest to date. The hypothesis of X17 boson has continuously received attentions by the data of anomalous 8Be, 4He, and 12C decays over time. The accumulation of data has shifted the most favored coupling strengths from $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ Feng:2016jff to $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ without isospin effects, and $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ with both isospin mixing and breaking effects considered Denton:2023gat . The persistent decay anomalies outlined above have motivated the testing of X17 hypothesis in the decays of heavy-flavor mesons such as $D^{*}$ and $B^{*}$ mesons with the assumption $\varepsilon_{b}=\varepsilon_{s}=\varepsilon_{d}$ and $\varepsilon_{c}=\varepsilon_{u}$ Castro:2021gdf . We have revisited this study by focusing on the comparisons of theoretical predictions (with X17 boson mediated effects included) on decay ratios $R_{D_{s}^{*+}}$ and $R_{D^{*0}}$ with experimental data CLEO:2011mla ; BESIII:2021vyq . For the case of $R_{D_{s}^{*+}}$, we have found that only the updated values of $\varepsilon_{u,d}$ are consistent with the measurement while the earlier extracted parameter values $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ predict a $R_{D_{s}^{*+}}$ in tension with the measurement (see Fig. 1). Our finding corrects an error by Ref. Castro:2021gdf which concluded that the prediction by $(\varepsilon_{u},\varepsilon_{d})=(\pm 3.7\times 10^{-3},\mp 7.4\times 10^{-3})$ is consistent with the measurement.
@@
To accommodate the measured $R_{D^{*0}}$, one requires a much larger $\left|\varepsilon_{u}\right|$ under the assumption $\varepsilon_{c}=\varepsilon_{u}$. On the other hand, raising the magnitude of $\varepsilon_{c}$ would affect $R_{D_{s}^{*+}}$ unless the magnitude of $\varepsilon_{s}$ is increased as well since $R_{D_{s}^{*+}}$ behaves approximately as $\left|\varepsilon_{c}+2\varepsilon_{s}\right|$ as indicated by Fig. 3 (left panel). Since $\varepsilon_{s}$ already departs from $\varepsilon_{d}$ in this case, we then also treat $\varepsilon_{u}$ and $\varepsilon_{c}$ as independent parameters in our analysis. To determine the favored ranges for $\varepsilon_{u}$, $\varepsilon_{c}$ and $\varepsilon_{s}$, we have performed a fitting to the data of $R_{D_{s}^{*+}}$, $R_{D^{*0}}$, $R_{\psi(2S)}$, and $R_{\phi}$. The results projected onto $(\varepsilon_{c},\varepsilon_{s})$, $(\varepsilon_{c},\varepsilon_{u})$, and $(\varepsilon_{s},\varepsilon_{u})$ planes are shown in Fig. 4. The best-fitted values for the coupling parameters are $\left|\varepsilon_{c}\right|=6.4\times 10^{-3}$, $\left|\varepsilon_{s}\right|=2.0\times 10^{-3}$, and two solutions for $\left|\varepsilon_{u}\right|$, which are $6.0\times 10^{-2}$ and $6.5\times 10^{-2}$, respectively.
@@
It is noteworthy that the value of $\left|\varepsilon_{u}\right|$ remains of the order $10^{-2}$ up to the $3\sigma$ range. Remarkably, this value is much larger than that determined from the anomalous decays of ${}^{8}\mathrm{Be},{}^{4}\mathrm{He}$, and ${}^{12}\mathrm{C}$, i.e., $\left|\varepsilon_{u}\right|=(0.5-0.9)\times 10^{-3}$ Denton:2023gat . Such a tension is caused by the large deviation of the measured $R_{D^{*0}}$ from the Standard Model prediction as shown in Table 2. This is also reflected by the right panel of Fig. 3, where both green bands significantly deviate from the line $\varepsilon_{u}=0$. In conclusion, we have studied effects of the proposed X17 boson to the decays of $D^{*0}$, $D^{*+}_{s}$, $\psi(2S)$, and $\phi$ mesons, which involve X17 boson couplings to both the first and second generations of quarks. The above-mentioned tension on the value of $\varepsilon_{u}$ deserves further scrutiny. We note that the decay $D^{*\pm}\to D^{\pm}e^{+}e^{-}$ remains to be measured. Furthermore the huge deviation of the measured $D^{*0}\to D^{0}e^{+}e^{-}$ branching ratio to its Standard Model predicted value awaits confirmations from updated measurements.